OpenStudy (anonymous):

this one is about factoring also...but a little bit nicer and harder\[p^x=y^4+4\]\(p\) is a prime...\(x,y\) are natural.

OpenStudy (anonymous):

What are we supposed to do?

OpenStudy (anonymous):

well...find all solutions

OpenStudy (anonymous):

this one is not mine

OpenStudy (anonymous):

it will be always odd because lhs is an odd no

OpenStudy (anonymous):

Oh.... got it can be both

OpenStudy (anonymous):

how...lhs is (prime no)^natural no....so odd ...right?

OpenStudy (anonymous):

p=2

OpenStudy (anonymous):

oh...yes...for p=2 it'll fail

OpenStudy (anonymous):

1...factor y^4+4 2...p^x=ab then a,b are some exponents of p

OpenStudy (anonymous):

y^4+4=(y^2-2y+2)(y^2+2y+2)

OpenStudy (anonymous):